y''+2y'+2y=2 y(0)=0,y'(0)=1


1
Expert's answer
2022-01-18T06:30:18-0500

Characteristic equation

"{k^2} + 2k + 2 = 0\\
D = 4 - 8 = - 4\\
{k_1} = \frac{{ - 2 - 2i}}{2} = - 1 - i\\
{k_2} = - 1 + i"

We obtain the general solution of the homogeneous equation

"{y_0} = {e^{ - x}}\left( {{C_1}\cos x + {C_2}\sin x} \right)"

We will seek a particular solution of the equation in the form

"Y = A \Rightarrow Y' = Y'' = 0"

then

"0 + 0 + 2A = 2 \Rightarrow A = 1"

Then

"Y = A,\,\,y = {y_0} + Y = {e^{ - x}}\left( {{C_1}\cos x + {C_2}\sin x} \right) + 1"


"y' = - {e^{ - x}}\left( {{C_1}\cos x + {C_2}\sin x} \right) + {e^{ - x}}\left( { - {C_1}\sin x + {C_2}\cos x} \right)"

"y(0) = 0,\,y'(0) = 1 \Rightarrow \left\{ {\begin{matrix}
{{C_1} +1= 0}\\
{ - {C_1} + {C_2} = 1}
\end{matrix}} \right. \Rightarrow {C_1} = -1,\,\,{C_2} = 0"

Answer: "y = - {e^{ - x}}\cos x + 1"